OpenStudy (anonymous):

Two trains leave the station at the same time, one heading west and the other east. The westbound train travels 20 miles per hour faster than the eastbound train. If the two trains are 320 miles apart after 2 hours, what is the rate of the westbound train?

OpenStudy (madhu.mukherjee.946):

The two trains are moving away from each other.

OpenStudy (madhu.mukherjee.946):

Two trains leave a railway station at the same time. The first train travels due west and the second due north. The first train travels 5km/hr faster than the second train. If after two hours, they are 50km apart, find the average speed of each train. let speed of 1st train=xkm/h let speed of 2nd train=x+5km/h t=2 hrs v=s/t a.t.q x+x+5=50/2 2x+5=25 2x=25­5=20 x=20/2=10km/h so, x=10km/hr x+5=15km/hr speed of 1st train=10km/hr speed of 2nd train=15km/hr

OpenStudy (madhu.mukherjee.946):

this is the sum which i had done ...probably this might help you!!

OpenStudy (zzr0ck3r):

West bound train has speed \(x+20\) where \(x\) is the speed of the other train. Since distance equals rate times time we get \[(x+20)t+x*t=320\] And we are told \(t=2\) so we have \[2(x+20)+2x=320\]

OpenStudy (madhu.mukherjee.946):

@mm41 do you understand

OpenStudy (madhu.mukherjee.946):

@ganeshie8 thanks

OpenStudy (zzr0ck3r):

West bound train has speed \(x+20\) where \(x\) is the speed of the other train. Since distance equals rate times time we get \[(x+20)t+x*t=320\] And we are told \(t=2\) so we have \[2(x+20)+2x=320\]

OpenStudy (anonymous):

so what is the answer

OpenStudy (madhu.mukherjee.946):

you put the values and you will get it

OpenStudy (zzr0ck3r):

or look at what I did and solve the last equation.

OpenStudy (madhu.mukherjee.946):

we can't give you direct answers here ,we can tell you the way and help you out but you got to get the answer on your own

OpenStudy (madhu.mukherjee.946):

@mm41 you calculate and give the answer we can check it out for you